If Sally can paint a house in 4 hours, and John can paint the same house in 6 hour, how long will it take for both of them to paint the house together?
Found 2 solutions by nycsharkman, oscargut:
Answer by nycsharkman(136)

If Sally can paint a house in 4 hours, and John can paint the same house in 6 hour, how long will it take for both of them to paint the house together?
Sally's time = 1/4
John's time = 1/6
together = 1/x
Solve for x.
1/4 + 1/6 = 1/x
5/12 = 1/x
Multiply both sides of the equation by the LCD, which is 12x
5/12 times 12x = 5x
1/x times 12x = 12
We now have a linear equation:
5x = 12
Divide both sides by 5 to find x.
x = 12/5
x = 2 hours and 40 minutes.
They should work together!

Answer by oscargut(2103)

Sally is doing 1/4 each hour John 1/6 each hour so together they work 1/4 + 1/6 = 10/24 = 5/12 in a hour.
5/12 -------> 1 hour
1 -------> x
then x=12/5 so the response is 12/5 hours = 2.4 hours = 2 hours and 24 minutes